# Strong Data Processing Properties of Rényi-divergences via Pinsker-type Inequalities

### v2

@@ IV Pinsker-type Inequalities for Rényi-divergence
For any $\alpha>1$, the bound in Theorem 8 is realized by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$. Hence, it is tight for this range of values. It is also tight for all $t\geq 0$ given that $\alpha=2$. In this case, it reduces to the joint-range-based bound between $\chi^{2}$-divergence and total variation distance, as e.g. presented in [3]. As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ hold for all $t>0$.
@@
An example of the the presented bound on the joint range for $\alpha=4$ is shown in Figure 1. Other values of $\alpha$ yield a visually similar result, with the discontinuity at $\nicefrac{{1}}{{\alpha}}$ moving closer to the origin as the value of $\alpha$ increases.
@@ V Strong Data Processing with Pinsker
$\Gamma_{\max}=\sup_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (33)
$\Gamma_{\min}=\inf_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (34)
@@
For a discrete channel $P_{Y|X}$, a set $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ and some $\alpha>1$, let $\eta\coloneqq\eta_{TV}(P_{Y|X})$. We have
$D_{f_{\alpha}}(P_{Y}||Q_{Y})\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(D_{f_{\alpha}}(P_{X}||Q_{X})).$ (35)
Proof:
Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$ and the inequality in Theorem 8. ∎
This result stresses the relation between $\Gamma_{\max}$ and $\Gamma_{\min}$ as formulated in (33) and (34) and the contraction behavior of a channel: If $\Gamma_{\max}-\Gamma_{\min}$ is small, the channel significantly contracts the $f_{\alpha}$-divergence. If the difference is large, contraction will be small. What remains is to bound $\Gamma_{\max}$ and $\Gamma_{\min}$ such that an operational bound is obtained. Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors. We present one such use-case specific bound in the following.
VI Application: Post-Processing for RLDP
In this section, we show how the above results can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [14, 25]. Consider the Markov chain $W-X-Y$, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, i.e.,
$\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}})\leq\varepsilon.$ (36)
Further, assume that the transition kernel $P_{Y|X}$ does not satisfy any finite RLDP guarantee, that is, the channel matrix contains zero elements. From the post-processing property of RLDP, it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$. However, it is reasonable to assume that there exist channels $P_{Y|X}$ containing zero elements that will result in an improved guarantee of the composed system $P_{Y|W}$ if the output of the RLDP $P_{X|W}$ is applied to their input. In fact, using the results derived above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
@@ VI Application: Post-Processing for RLDP
$\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}\biggl{(}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})\biggr{)},$ (39)
where $\eta\coloneqq\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})$ is again a shorthand for the contraction coefficient of the total variation distance. Since $P_{X|W}$ satisfies $(\varepsilon,\alpha)$-RLDP, we can therefore bound the RLDP guarantee of the entire system by applying (4) as
$\varepsilon^{*}\coloneqq\text{RLDP}(P_{Y|W})\leq\varphi(\varepsilon_{f},\alpha),$ (40)
where $\varepsilon_{f}\coloneqq\max_{w,w^{\prime}}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})$ is the $f_{\alpha}$-divergence corresponding to the $\varepsilon$-valued Rényi-divergence,
$\varphi(\varepsilon_{f},\alpha)\coloneqq\frac{1}{\alpha-1}\log\biggl{(}\eta_{TV}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(\varepsilon_{f})+1\biggr{)}.$ (41)
Further, in this setup, the set of input distributions $\mathcal{P}$ is just the set of rows of the mechanism $P_{X|W}$, and each of these distributions is induced by one of the outcomes $W=w$. Hence, the quantities $\Gamma_{\max}$ and $\Gamma_{\min}$ reduce to
$\Gamma_{\max}=\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)},$ (42)
$\Gamma_{\min}=\min_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}.$ (43)
While these quantities can be computed directly if both channels are perfectly known, it is useful to obtain a bound that only depends on the privacy guarantees of the channels. The following proposition provides such a bound, where the privacy of the second, non-RLDP channel is quantified by its pointwise maximal leakage (PML) [21]. The proposition shows one way in which the increased flexibility of leakage assessment with PML can complement (L)DP guarantees.
Proposition 12.
If $P_{X|W}$ is a $N\times N$ randomized response mechanisms satisfying $\varepsilon$-LDP according to [31], then
$\Gamma_{\max}\leq\biggl{(}1-\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (44)
$\Gamma_{\min}\geq\biggl{(}1+\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (45)
where $\ell_{w}(X\to y)$ denotes the PML from $X$ to an outcome $Y=y$ under the input distribution induced by $W=w$, i.e.,
$\ell_{w}(X\to y)\coloneqq\log\frac{\max_{x\in\mathcal{X}}P_{Y|X=x}(y)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}.$ (46)
Further, equality is achieved in both bounds if $P_{Y|X}$ has a zero entry in each column, i.e., if $\min_{x}P_{Y|X=x}(y)=0$ $\forall y\in\mathcal{Y}$.
Proof:
See Appendix C. ∎
The following example illustrates this bound.
Figure 2: RLDP and corresponding bound in Example 13.
@@
Let $P_{X|W}$ be a $5\times 5$ randomized response mechanism in [31] satisfying $\log 6$-LDP, and $P_{Y|X}$ given by
$P_{Y|X}=\begin{bmatrix}\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0&0\\ 0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0\\ 0&0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}\\ \nicefrac{{1}}{{2}}&0&0&0&\nicefrac{{1}}{{2}}\end{bmatrix}.$ (47)
[…]
Appendix A Proof of Proposition 3
For $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$, we have

### v3

@@ IV Pinsker-type Inequalities for $f_{\alpha}$-divergence
For any $\alpha>1$, the bound in Theorem 8 is realized by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$. Hence, it is tight for this range of values. It is also tight for all $t\geq 0$ given that $\alpha=2$. In this case, it reduces to the joint-range-based bound between $\chi^{2}$-divergence and total variation distance, as e.g. presented in [2]. As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ hold for all $t>0$. Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=0.$ (40)
In the range $t\in(0,\nicefrac{{1}}{{\alpha}})$, this equation has a real solution $p^{*}$ that satisfies $p^{*}\in[0,1]$. Due to the convexity of $\psi_{\alpha}$ in $p$, a (tight) lower bound is then given by $g_{\alpha}(t)=\psi_{\alpha}(p^{*},t)$. However, it remains unclear if $p^{*}$ can be expressed in closed form. Due to this complication, the result above utilizes more rudimentary bounds in this case. Further, we invoke standard Pinsker’s inequality for relative entropy to derive the first case of (39). This bound is not the best we can do for this case. More specifically, the optimal Pinsker’s inequality has been found in [26] in implicit form. We rely on the more rudimentary standard Pinsker bound here for simplicity.
@@
An example of the presented bound on the joint range for $\alpha=4$ is shown in Figure 1. Other values of $\alpha$ yield a visually similar result, with the discontinuity at $\nicefrac{{1}}{{\alpha}}$ moving closer to the origin as the value of $\alpha$ increases.
@@ V Strong Data Processing with Pinsker
$\Gamma_{\max}(Y)=\sup_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)},$ (42)
$\Gamma_{\min}(Y)=\inf_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)},$ (43)
@@
For a discrete channel $P_{Y|X}$, distributions $P_{X},Q_{X}\in\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$, $P_{Y}=P_{Y|X}\circ P_{X}$, $Q_{Y}=P_{Y|X}\circ Q_{X}$ and some $\alpha>1$, we have
$D_{f_{\alpha}}$ $(P_{Y}||Q_{Y})$ (44)
$\leq\eta_{TV}(P_{Y|X},\mathcal{P})R_{\alpha}(\Gamma_{\max}(Y),\Gamma_{\min}(Y))$ (45)
$\qquad\qquad\qquad\qquad\qquad\times g_{\alpha}^{-1}\Big{(}D_{f_{\alpha}}(P_{X}||Q_{X})\Big{)}.$ (46)
Proof.
Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$, the inequality in Theorem 8, the inverse given in Remark 10 and since by definition, $\text{TV}(P_{Y}||Q_{Y})\leq\eta_{TV}(P_{Y|X},\mathcal{P})\text{TV}(P_{X}||Q_{X})$ if $P_{X},Q_{X}\in\mathcal{P}$. ∎
This result stresses the relation between $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ as formulated in (42) and (43) and the contraction behavior of a channel: If $\Gamma_{\max}(Y)-\Gamma_{\min}(Y)$ is small, the channel significantly contracts the $f_{\alpha}$-divergence. If the difference is large, contraction will be small. What remains is to bound $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ such that an operational bound on contraction is obtained. Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors. One such bound appears in the application of these results to privacy amplification, which we will present next.
VI Privacy Amplification Bounds for RLDP
In this section, we show how the above technical tools can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [27, 5]. Consider the Markov chain WXYW-X-Y, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, [^1] The notion of Rényi local differential privacy as a local adaption of Rényi differential privacy [27] was recently suggested by Gilani et al. [5], where the authors refer to it as local Rényi differential privacy (LRDP). Note that the class of discrete mechanisms satisfying any (finite) RLDP guarantee for any $\alpha$ is equivalent to the class of discrete mechanisms satisfying some (finite) LDP guarantee. Specifically, this class only includes mechanisms with strictly positive entries in the mechanism matrix. i.e.,
$\text{RLDP}^{(\alpha)}$ $(P_{X|W})$ (47)
$\coloneqq\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}})\leq\varepsilon.$ (48)
Further, assume that the channel $P_{Y|X}$ does not satisfy any finite RLDP guarantee. From the post-processing property of RLDP (which is just the DPI for Rényi-divergence), it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$. However, it is reasonable to assume that there are certain $P_{Y|X}$ with zero-valued elements that improve the privacy guarantee of $P_{Y|W}$ when cascaded with the RLDP channel $P_{X|W}$. Using our results above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
@@ VI Privacy Amplification Bounds for RLDP
$\qquad\leq\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})R_{\alpha}(\Gamma_{\max}(Y),\Gamma_{\min}(Y))$ (51)
$\qquad\qquad\qquad\quad\times g_{\alpha}^{-1}\Big{(}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})\Big{)}.$ (52)
Since $P_{X|W}$ satisfies $(\varepsilon,\alpha)$-RLDP, we can apply (4) to bound the RLDP guarantee of the entire system as
$\varepsilon^{*}\coloneqq\text{RLDP}^{(\alpha)}(P_{Y|W})\leq\varphi(\varepsilon_{f},\alpha),$ (53)
where $\varepsilon_{f}\coloneqq\max_{w,w^{\prime}}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})$ is the $f_{\alpha}$-divergence corresponding to the $\varepsilon$-valued Rényi-divergence, and we define $\varphi:[0,\infty)\times(1,\infty)\to\mathbb{R}$ as
$\varphi$ $(\varepsilon_{f},\alpha)$ (54)
$\coloneqq\frac{1}{\alpha-1}\log\biggl{(}\eta_{TV}R_{\alpha}(\Gamma_{\max}(Y),\Gamma_{\min}(Y))g_{\alpha}^{-1}(\varepsilon_{f})+1\biggr{)}.$ (55)
Further, in this setup, the set of input distributions $\mathcal{P}$ to the channel $P_{Y|X}$ is just the set of rows of the RLDP mechanism $P_{X|W}$, and each of these distributions is induced by one of the outcomes $W=w$. Hence, the quantities $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ reduce to
$\Gamma_{\max}(Y)=\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)},$ (56)
$\Gamma_{\min}(Y)=\min_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}.$ (57)
The following example illustrates the above results.
(a) $N=5$, $\varepsilon=\log 2$, $P_{Y|X}$
(b) $N=5$, $\varepsilon=\log 6$, $P_{Y|X}$
(c) $N=20$, $\varepsilon=\log 10$, $P_{Y|X}$
(d) $N=100$, $\varepsilon=\log 10$, $Q_{Y|X}$
Figure 2: RLDP guarantees and the corresponding bounds in Example 12.
@@
Let $P^{(N)}_{X|W}$ be the $N\times N$ randomized response mechanism in [28] satisfying $\varepsilon$-LDP, that is, let
$P^{(N)}_{X|W=w}(x)=\begin{cases}\frac{e^{\varepsilon}}{N+e^{\varepsilon}-1},&\text{ if }x=w\\ \frac{1}{N+e^{\varepsilon}-1},&\text{ otherwise}.\end{cases}$ (58)
We examine the privacy amplification of the channels $P_{Y|X}$ and $Q_{Y|X}$ defined for some $N$ and $i,j\in[N]$ by
$P^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\nicefrac{{1}}{{2}},&\text{ if }i=j,\\ \nicefrac{{1}}{{2}},&\text{ if }i=j+1\mod N,\\ 0,&\text{otherwise}\end{cases}$ (59)
and for even values of $N$,
$Q^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\frac{2}{N},&\text{if }1\leq i,j\leq\frac{N}{2}\\ \frac{2}{N},&\text{if }\frac{N}{2}<i,j\leq N\\ 0,&\text{otherwise}\end{cases}$ (60)
In this setup, the successive application of both mechanisms to the private data $W$ will increase randomization, and hence it is reasonable to expect an improved privacy guarantee of the complete system compared to $P^{(N)}_{X|W}$ alone. We can numerically determine the bound in (53) assuming the upper bound $\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. The resulting RLDP values are shown in Figure 2 alongside the RLDP guarantee of $P^{(N)}_{X|W}$ (the best bound previously known) and the true RLDP guarantee of the composition $P^{(N)}_{X|W}\circ P^{(N)}_{Y|X}$ or $P^{(N)}_{Y|X}\circ P_{X|W}$. As expected, the bound tightens as $\alpha$ grows, since both of the presented Pinsker-type bounds are asymptotically tight. Computing the values of $\Gamma_{\max}$ and $\Gamma_{\min}$ in this setup, we find that given the channel $P^{(N)}_{Y|X}$,
$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{e^{\varepsilon}+1}{2}.$ (61)
On the other hand, given $Q^{(N)}_{Y|X}$, we have
$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{N+2e^{\varepsilon}-2}{N}.$ (62)
That is, the contraction behavior of the channels $P^{(N)}_{Y|X}$—and hence its privacy amplification—is independent of the alphabet size $N$. On the other hand, the amplification of the channels with $Q^{(N)}_{Y|X}$ crucially depends on the alphabet size. Interestingly, as long as $\varepsilon<\infty$, we see that asymptotically, $\lim_{N\to\infty}\Gamma_{\max}(Y)=\lim_{N\to\infty}\Gamma_{\min}(Y)=1$, that is, the privacy amplification of the channels $Q^{(N)}_{Y|X}$ becomes independent of the privacy guarantee of the system $P^{(N)}_{X|W}$. In particular, post-processing a randomized response mechanism with any finite LDP guarantee with a channel of a structure similar to $Q^{(N)}_{Y|X}$ will result in a system with perfect privacy as $N\to\infty$.
Example 12 shows that even sparse channel matrices can lead to significant privacy amplification if applied as a post-processing step to an LDP system. Moreover, as long as the number of non-zero elements in the channel matrix grows linearly with $N$, the privacy guarantee of the system will approach prefect privacy as the alphabet size grows. While these insights are so far limited to randomized response mechanisms, they are interesting for two main reasons:
First, the results motivate us to further investigate arbitrary systems that apply some form of randomization in terms of their effect on LDP guarantees; it has been previously observed that local differential privacy constraints on a system can be severely limiting, especially in large-scale learning applications [29, 30]. In this context, the results presented in this paper are encouraging. Specifically, if some form of LDP processing is done to the data, even a learning system without a formal LDP guarantee on its own might be able to significantly increase the privacy guarantee with respect to the data used to train it.
Second, the results are promising for addressing the common issue of large values of $\varepsilon$ in $\varepsilon$-LDP guarantees in practical systems [31]: Instead of fine tuning the practical application itself to improve its privacy guarantee (which might be intractable), practitioners might be able to apply simple post-processing steps to an application output in order to significantly improve the systems privacy guarantee, while retaining much of the original systems performance.
Future work in this direction should focus on more advanced tools and methods for bounding the values $\Gamma_{\max}$ and $\Gamma_{\min}$ for various (practical) systems. Obtaining such bounds paves the way for easy plug-and-play solutions for improving privacy guarantees of weak LDP systems at a minimal loss of utility.
Appendix A Proof of Theorem 8
where $d_{f_{\alpha}}(p||q)$ denotes the binary $f_{\alpha}$-divergence
Define
From the definition of the total variation, we get $\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q)=\min_{p\in[0,1-t]}\{\psi_{\alpha}(p,t)\}$. Taking derivatives with respect to $p$ yields
and for any $\alpha>1$ and all $t<1,\,p\leq 1-t$,
from which we can conclude that $\psi_{\alpha}(p,t)$ is convex in $p$.
with equality for $t=\nicefrac{{1}}{{\alpha}}$ if $p=0$. Otherwise strict inequality holds. Together with the convexity of $\psi_{\alpha}$ in $p$, this implies that the maximum value of $\psi_{\alpha}$ for all $t\geq\nicefrac{{1}}{{\alpha}}$ will be achieved at $p=0$. At this point, we have $\psi_{\alpha}(0,t)=(1-t)^{1-\alpha}$.
$=\biggl{(}1+\frac{t^{2}}{(p+t)(1-p-t)}\biggr{)}^{\alpha-1}-1$
$\geq(1+4t^{2})^{\alpha-1}-1.$
Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [4], so is $D_{f_{\alpha}}$. Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$), $\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$, will continue to hold for $\alpha>1$. Utilizing (4), this yields
Appendix B Proof of Proposition 3
For some fixed $P_{Y|X}$, let $P_{Y}=P_{Y|X}\circ P_{X}$ and $Q_{Y}=P_{Y|X}\circ Q_{X}$. For $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$, we have